The Information Systems Audit and Control Association surveyed office workers to learn about the anticipated usage of office computers for personal holiday shopping (USA Today, November 11, 2009). Assume that the number of hours a worker spends doing holiday shopping on an office computer follows an exponential distribution. a. The study reported that there is a .53 probability that a worker uses the office computer for holiday shopping 5 hours or less. Round your answers to four decimal places. Is the meantime spent using an office computer for holiday shopping closest to 5.8, 6.2, 6.6, or 7 hours?

*it is 6.6- I got that correct*

I need help with the following probabilities:

μ Probability

5.8 ----------

6.2 -----------

6.6 -----------

7.0 ------------
Answer #1

for exponential distribution: P(X=x) =1-e-x/ μ

μ Probability
5.8 =1-exp(-5/5.8)=0.5777
6.2 =1-exp(-5/6.2)=0.5536
6.6 =1-exp(-5/6.6)=0.5312
7 =1-exp(-5/7)=0.5105